OpenStudy (anonymous):

A system of linear inequalities is shown below: y ≥ -2x - 2 y + x ≥ 0 Describe the steps to graph the solution set to the system of inequalities. (10 points)

OpenStudy (anonymous):

@texaschic101 could you please please help me

OpenStudy (anonymous):

@jim_thompson5910 help

OpenStudy (texaschic101):

do you know how to graph equations ? You can pretty much do the same thing with inequalities except for a few things. First put the inequality in y = mx + b form where m is the slope and b is the y intercept. y >= -2x + 2 -- disregard the inequality sign and sub in a 0 (for the moment) y = -2x + 2 --- so the slope is -2 and the y intercept is (0,2) the x intercept can be found by subbing in 0 for y y = -2x + 2 0 = -2x + 2 2x = 2 x = 1.....so the x intercept is (1,0) Now we know the intercepts are (0,2) and (1,0) Now put back in your inequality sign .. y >= -2x + 2.. if it has an equal sign, the line is solid...if it is greater then, shading occurs above the line. So plot your intercepts....connecting them with a solid line...then shade above the line. That is one line...with this info, do you think you can graph the other line by doing the exact same thing I did for this line ?

OpenStudy (texaschic101):

oops...disregard the inequality sign and sub in an = sign

OpenStudy (anonymous):

i think i could do the other line thanks alot :) if i get stuck on something could i send you a message?

OpenStudy (texaschic101):

of course you can

OpenStudy (anonymous):

k thanks :)